# Proton Radiation Damage and Annealing of COSI p-type Cross-strip HPGe Detectors

### v1

@@ Abstract
In order to understand the effects of a space radiation environment on cross-strip germanium detectors, we investigated the effects of high-energy proton damage on a COSI detector and the capabilities of high-temperature annealing in repairing detector spectral resolution. We irradiated a COSI-balloon cross-strip high-purity germanium (HPGe) detector with a high-energy proton fluence corresponding to 10 years in a space radiation environment. We repaired the resulting degradation in spectral resolution within 16 of its preradiation value through a series of high-temperature anneals. We characterize the repair of charge traps with time spent under high-temperature anneal to inform an annealing procedure for long-term maintenance of COSI’s spectral resolution.
@@ 1.1 Background
The Compton Spectrometer and Imager (COSI) is a NASA Small Explorer (SMEX) satellite mission currently under development for launch in 2027 [1]. COSI is a wide-field gamma-ray telescope designed to provide spectroscopy, imaging, and polarimetry of astrophysical sources in the bandpass 0.2-5 MeV. The heart of the COSI instrument is an array of 16 cross-strip germanium detectors (GeDs) providing excellent energy resolution to achieve two of the primary mission goals. First, to uncover the origin of Galactic positrons through imaging spectroscopy of the annihilation emission at 0.511 MeV. And second, to reveal Galactic element formation through study of gamma-ray emission lines produced through radioactive decay of recently synthesized elements. To fulfill these science goals, COSI must maintain excellent spectral performance throughout its primary mission.
@@
The spectral resolution of GeDs in space degrade due to increased charge trapping induced by radiation damage from ionizing particles, most significantly high-energy protons. High-energy protons damage the crystal lattice structure of the detectors and produce negatively ionized deformations which trap holes as they drift in the applied bias voltage towards the signal electrodes [2]. Trapping degrades spectral resolution of the hole-collection signal, characterized by the broadening of the photpeaks, as measured by the full width at half maximum (FWHM), and increased low-energy tailing of spectral lines. This degrades the ability to determine the energy of the incoming gamma-rays.
@@
Similar radiation damage effects have been observed by germanium spectrometers INTEGRAL/SPI [3] and RHESSI [4] during their time in orbit. Both instrument teams implemented a series of high-temperature anneals at 100∘C to repair the radiation damage and restore the spectral resolution. However, annealing at these temperatures in orbit requires careful planning and design to avoid permanent damage to any instrument. The SPI telescope is composed of 19 coaxial GeDs launched into a highly eccentric orbit (perigee of 9000 km and apogee of 154600 km). After noting a significant degradation in energy resolution after 2 years in orbit, SPI began completing periodic high-temperature anneals to repair spectral resolution and prevent permanent damage from radiation [5]. SPI currently anneals at 100∘C for 14 days every 6 months, observing an improvement of 10-15 to their spectral resolution between each anneal [6]. However, they suffered the loss of 2 GeDs [5]. In low earth orbit (LEO) at 38∘ inclination, RHESSI similarly experienced significant losses in resolution which warranted an annealing procedure. Nine segmented coaxial n-type GeDs underwent 5 anneals at 100∘C over 9 years of operation, each anneal lasting for approximately 1 month [7]. Although the length of each anneal was minimized to mitigate the risk of fusing the spectrometer’s segmented germanium crystals, one of RHESSI’s detectors became unsegmented after multiple anneals [8].
@@
Neutron radiation induced damage to GeDs has been studied extensively [2, 12, 13, 14]. However, relatively few studies of proton irradiation on detector performance provide insight for this work. Initial study of proton radiation was carried out on planar p-type GeDs with 6 GeV protons providing fluences ranging (0.73 - 5.7) p+/cm2 [15]. The effects of irradiation included prominent low-energy tailing and an increase in FWHM which was approximately linear with fluence. The measurements from these studies concluded that protons induce damage at lower fluences than neutron irradiations. These observations, paired with a shorter timescale for annealing repair of proton damage, suggest that the damaging mechanism for protons is distinct from that of neutrons. 1.5 GeV proton irradiation of p- and n-type coaxial GeDs with net fluences on each detector within the range of p+/cm2 likewise produced significant increase in FWHM and tailing at low energies [16]. Testing three detectors in parallel, while holding each at a different temperature during irradiation, revealed a greater degradation in spectral resolution at higher detector temperatures.
Our objectives for this work are to characterize the effects of radiation damage on COSI’s spectral performance over the primary and extended missions, and to formulate and test an annealing procedure to repair said damage and restore COSI’s spectral resolution to its baseline value. Through these damage and annealing procedures, we track the spectral resolution by measuring the FWHM of the spectral line from decaying isotope Cesium 137 (Cs-137), emitting gamma-rays at 0.662 MeV. We utilize numerical charge transport simulations, including a model of trapping [17] to relate the density of charge traps within the detector volume to the resulting degradation in spectral resolution, and ultimately track how trapping defects decrease with time spent annealing at high temperatures.
@@ 2.1 Primary and Secondary Traps
Secondary defects form as the result of interstitial motion which takes place when the detectors are cycled to room temperature. Fourches et al. (1991b) used deep-level transient spectroscopy (DLTS) to track signatures of traps, and concluded that secondary damage is caused by the migration of individual vacancies out of the primary disordered regions, potentially combining to form divacancy sites. These sites also act as hole traps. Confirmed in previous works [19, 14] and in our experience temperature cycling this detector, secondary defects do not form when undamaged germanium is cycled to room temperature, only when primary defects are already present. The formation of secondary defects is important to understanding the potential impacts of radiation damage as previous studies observe a more significant degradation in resolution from secondary defects than primary defects [12, 14].
@@ 2.2 Trapping Parameters
As derived in Boggs Pike (2023), the total path length is the vector combination of the net drift displacement and random thermal motions around the path of travel. Given the carrier drift velocity, , and random thermal velocity, , (Figure 1), the total distance traveled over time is given by:
We characterize trapping via the product , termed the ‘trapping product’ and measured in units of cm. Under this definition, no charge trapping would correspond to an infinite trapping product. The lower the trapping product, the more trapping that occurs for a given path length . Neither nor can be measured directly from our detector resolution measurements. However, using simulations of our detectors for various trapping products, we can derive an empirical relation between and the resulting spectral resolution, which we discuss further in Section .
@@ 3 Methods
For this study, we used a spare p-type high purity Germanium (HPGe) cross-strip detector from the COSI-Balloon mission [20]. The detector has dimensions of 8 cm x 8 cm x 1.5 cm and 37 electrode strips of 2-mm pitch on each face. This detector depletes at 400V, and we operate it at an applied bias of +600V within an evacuated cryostat cooled by liquid nitrogen to 80K. A gamma-ray interacting with the germanium crystal generates a cloud of electron-hole pairs from the material. Freed charges drift in the applied electric field, with the hole charge cloud collected on the LV (ground) electrode strip and the electron charge cloud on the HV (high voltage) electrode strip, as pictured in Figure 1. Each electrode is connected to readout electronics which process the signals; These are the exact readout electronics used on the COSI-Balloon payload as described in Beechert et al (2020). Events are then binned into energy spectra for analysis. To calibrate the detector and measure the spectral resolution prior to any radiation or annealing, we collected photopeak events from gamma-ray sources across a range of energies in the MeV band.
In this work, we focus on spectral measurements of the 0.662 MeV line from Cs-137. We measure the spectral resolution of the electron-collecting (AC) and hole-collecting (DC) strips separately, which enables us to distinguish between the effects of electron and hole trapping.
@@ 3.1 Radiation Damage
In order to replicate the proton fluence we might expect over COSI’s mission, we used the proton synchrotron at the James M Slater MD Proton Treatment and Research Center at Loma Linda University Medical Center. We irradiated our detector in two rounds with protons with an average energy of 150 MeV incident on the HV face of the detector. During the first round, the detector received a fluence of p+/cm2. Then, after measuring the spectral resolution of the detector, we delivered another p+/cm2, resulting in a net fluence of p+/cm2. During irradiation, the detector was kept under vacuum at operating temperature ( K). Our measurements of the spectral resolution after each round of radiation revealed a significant broadening in photopeaks of hole-collecting strips, characteristic of increased hole trapping due to radiation damage. After completing our post-irradiation measurements, we cycled the detector to room temperature for 336 hours while keeping it under vacuum.
@@ 3.2 Annealing
Anneal | Temperature | Time Annealing |
(∘C) | (Hours) | |
Room Temp | 336 | |
1 | 80 | 24 |
2 | 80 | 48 |
3 | 100 | 24 |
4 | 100 | 48 |
5 | 100 | 96 |
6 | 100 | 192 |
7 | 100 | 192 |
After each annealing procedure we returned the detector to its operating temperature and collected data from a Cs-137 source, tracking the spectral resolution through the resulting FWHM of the 0.662 MeV line. Radiation damage and annealing at room temperature produced non-gaussian line profiles including low-energy tailing, spectral line shifts, and bimodal distributions. Therefore we chose to fit the Cs-137 peaks with a univariate spline interpolation in order to obtain a FWHM measurement. We also fit the preradiation datasets using this method to maintain consistent FWHM measurements.
To probe the effectiveness of annealing procedures at different temperatures, we completed anneals at C and C. The 5 anneals at C were intended to characterize the progress of the spectral resolution back towards its preradiation value.
@@ 3.3 Numerical Simulations
In parallel to our measurements characterizing the effects of radiation and annealing on the spectral performance, we used numerical simulations to model the photopeak spectrum for a range of trapping products encompassing the anticipated levels of radiation damage. Our numerical charge transport simulations, including the model for charge trapping, are described and tested in [17].
The measured FWHM of the photopeak has two components: FWHMd attributed to the trap density caused by proton damage, and FWHM0 attributed to a combination of intrinsic trap density, statistical Fano noise, and noise from electronic readout. These components add in quadrature:
For hole-collecting strips, we equated FWHM0 to the measured preradiation FWHMh ( 2.67 keV). We tabulated the FWHM of the Cs-137 photopeak produced by our numerical simulations against their respective trapping products in order to determine a relationship between the two parameters. We interpolate this relation to pair the measured FWHM produced at each stage of damage and annealing with the corresponding hole trapping products. These FWHM values and corresponding trapping products are listed in Table 2.
@@ 4 Results
FWHMh | FWHMd,h | ||
(keV) | (keV) | (cm) | |
Preradiation | 2.67 | – | 18601 1 footnotemark: |
Radiation 1 | 17.99 | 17.78 | 104 |
Radiation 2 | 50.36 | 50.29 | 41 |
Room Temp | 10 | ||
Anneal 1 | 38.11 | 30.01 | 49 |
Anneal 2 | 21.11 | 21.06 | 92 |
Anneal 3 | 19.28 | 19.09 | 97 |
Anneal 4 | 17.03 | 16.81 | 114 |
Anneal 5 | 7.44 | 6.94 | 295 |
Anneal 6 | 3.09 | 1.55 | 893 |
Anneal 7 | 3.13 | 1.63 | 873 |
@@ 4.1 Preradiation
We measured the FWHM of the Cs-137 photopeak for hole and electron-collecting strips of 2.67 keV and 4.27 keV respectively, prior to any radiation damage or annealing. The broader resolution on the electron-collecting strips is due to AC-coupling and intrinsic electron trapping [21]. The photopeak for hole-collecting strips pictured in (Figure 4).
@@ 4.2 Post Radiation
After the first round of proton irradiation, we measure a peak with a 17.99 keV FWHM for hole-collecting strips – increased by a factor of six from preradiation. After the second round of irradiation, our measured FWHM increased to 50.36 keV, over 18 times its preradiation value. In terms of fluence, this reflects an increase of roughly 8.89 keV per p+/cm2 after the first round of radiation and 10.15 keV per p+/cm2 after the second round of radiation. These measurements demonstrate a roughly linear relationship between detector resolution and net proton fluence. This trend is confirmed in Fourches (1991b) and Fourches (1991a) for p-type GeDs, and in Kandel (1999) for SPI’s n-type GeDs, all of which underwent neutron irradiation. Likewise, this is seen in in Pehl (1978) for p-type GeDs which underwent proton irradiation. We observe widened photopeaks and low-energy tailing (Figure 2) characteristic of radiation damage.

### v2

@@ Abstract
In order to understand the effects of a space radiation environment on cross-strip germanium detectors, we investigated the effects of high-energy proton damage on a COSI detector and the capabilities of high-temperature annealing in repairing detector spectral resolution. We irradiated a COSI-balloon cross-strip high-purity germanium (HPGe) detector with 150 MeV protons resulting in a net fluence of p+/cm2 and corresponding to 10 years in COSI’s space radiation environment. We repaired the resulting degradation in spectral resolution through a series of high-temperature anneals to obtain a final FWHM of 4.08 keV, within 37 of its preradiation value (2.98 keV FWHM). We characterized the repair of charge traps with time spent under high-temperature anneal to inform an annealing procedure for long-term maintenance of COSI’s spectral resolution.
@@ 1.1 Background
The Compton Spectrometer and Imager (COSI) is a NASA Small Explorer (SMEX) satellite mission currently under development for launch in 2027 [1]. COSI is a wide-field gamma-ray telescope designed to provide spectroscopy, imaging, and polarimetry of astrophysical sources in the bandpass 0.2-5 MeV. The heart of the COSI instrument is an array of 16 cross-strip germanium detectors (GeDs) providing excellent energy resolution (0.5% at 0.662 MeV for single-strip events in Beechert et al. [2]) to aid in two of the primary mission goals. First, to uncover the origin of Galactic positrons through imaging spectroscopy of the annihilation emission at 0.511 MeV. And second, to reveal Galactic element formation through study of gamma-ray emission lines produced through radioactive decay of recently synthesized elements. To fulfill these science goals, COSI must maintain its spectral performance throughout its primary mission.
@@
The spectral resolution of GeDs in space degrade due to increased charge trapping induced by radiation damage from ionizing particles, most significantly high-energy protons. High-energy protons damage the crystal lattice structure of the detectors and produce negatively ionized deformations which trap holes as they drift in the applied bias voltage towards the signal electrodes [3]. Trapping degrades spectral resolution of the hole-collection signal, characterized by the broadening of the photopeaks, as measured by the full width at half maximum (FWHM), and increased low-energy tailing of spectral lines. This degrades the ability to determine the energy of the incoming gamma-rays.
@@
Similar radiation damage effects have been observed by germanium spectrometers INTEGRAL/SPI [4] and RHESSI [5] during their time in orbit. Both instrument teams implemented a series of high-temperature anneals at 100∘C to repair the radiation damage and restore the spectral resolution. However, annealing at these temperatures in orbit requires careful planning and design to avoid permanent damage to any instrument. The SPI telescope is composed of 19 coaxial GeDs launched into a highly eccentric orbit (perigee of 9000 km and apogee of 154600 km). After noting a significant degradation in energy resolution after 2 years in orbit, SPI began completing periodic high-temperature anneals to repair spectral resolution and prevent permanent damage from radiation [6]. SPI currently anneals at 100∘C for 14 days every 6 months, observing an improvement of within 10-15 to their spectral resolution between each anneal [7]. However, they suffered the loss of 2 GeDs [6]. In low earth orbit (LEO) at 38∘ inclination, RHESSI similarly experienced significant losses in resolution which warranted an annealing procedure. Nine segmented coaxial n-type GeDs underwent 5 anneals at 100∘C over 9 years of operation, each anneal lasting for approximately 1 week [8]. Although the length of each anneal was minimized to mitigate the risk of fusing the spectrometer’s segmented germanium crystals at the lithium-drifted contacts, multiple RHESSI detectors became permanently unsegmented after the fourth anneal [9].
@@
Neutron radiation induced damage to GeDs has been studied extensively [3, 13, 14, 15]. However, relatively few studies of proton irradiation on detector performance provide insight for this work. Initial study of proton radiation was carried out on planar p-type GeDs with 6 GeV protons providing fluences ranging (0.73 - 5.7) p+/cm2 [16]. The effects of proton irradiation included prominent low-energy tailing and an increase in FWHM which was approximately linear with fluence. The measurements from these studies concluded that GeV protons induce damage at lower fluences than irradiation from MeV neutrons. Furthermore, sharp peaked distributions post proton damage suggest that the damaging mechanism for protons is distinct from that of neutrons. 1.5 GeV proton irradiation of p- and n-type coaxial GeDs with net fluences on each detector within the range of p+/cm2 likewise produced significant increase in FWHM and tailing at low energies [17]. Testing three detectors in parallel, while holding each at a different temperature during irradiation, revealed a greater degradation in spectral resolution at higher detector temperatures within a range of 90 K to 130 K.
Our objectives for this work are to characterize the effects of radiation damage on COSI’s spectral performance over the primary and extended missions, and to formulate and test an annealing procedure to repair said damage and restore COSI’s spectral resolution to its baseline value. Through these damage and annealing procedures, we track the spectral resolution by measuring the effective FWHM of the spectral line from decaying isotope Cesium 137 (Cs-137), emitting gamma-rays at 0.662 MeV. We utilize numerical charge transport simulations, including a model of trapping [18] to relate the density of charge traps within the detector volume to the resulting degradation in spectral resolution, and ultimately track how trapping defects decrease with time spent annealing at high temperatures.
@@ 2.1 Primary and Secondary Traps
Secondary defects form as the result of interstitial motion which takes place when the detectors are cycled to room temperature. Fourches et al. [13] used deep-level transient spectroscopy (DLTS) to track signatures of traps, and concluded that secondary damage is caused by the migration of individual vacancies out of the primary disordered regions, potentially combining to form divacancy sites. These sites also act as hole traps. Confirmed in previous works [20, 15] and in our experience temperature cycling this detector, secondary defects do not form when undamaged germanium is cycled to room temperature, only when primary defects are already present. The formation of secondary defects is important to understanding the potential impacts of radiation damage as previous studies observe a more significant degradation in resolution from secondary defects than primary defects [13, 15].
@@ 2.2 Trapping Parameters
As derived in Boggs et al. [18], the total path length is the vector combination of the net drift displacement and random thermal motions around the path of travel. Given the carrier drift velocity, , and random thermal velocity, the total distance traveled over time is given by:
We characterize trapping via the product , termed the ‘trapping product’ and measured in units of cm. Under this definition, no charge trapping would correspond to an infinite trapping product. The lower the trapping product, the more trapping that occurs for a given path length . Neither nor can be measured directly from our detector resolution measurements. However, using simulations of our detectors for various trapping products, we can derive an empirical relation (Figure 1) between and the resulting spectral resolution, which we discuss further in Section 3.3.
@@ 3 Methods
For this study, we used a spare p-type high purity Germanium (HPGe) cross-strip detector from the COSI-Balloon mission [2]. The detector, shown in Figure 2, has dimensions of 8 cm x 8 cm x 1.5 cm and 37 electrode strips of 2-mm pitch on each face. This detector depletes at 400V, and we operate it at an applied bias of +600V within an evacuated cryostat cooled by liquid nitrogen to 80K. A gamma-ray interacting with the germanium crystal generates a cloud of electron-hole pairs from the material. Freed charges drift in the applied electric field, with the hole charge cloud collected on the LV (ground) electrode strip and the electron charge cloud on the HV (high voltage) electrode strip, as pictured in Figure 3. Each electrode is connected to readout electronics which process the HV and LV signals separately; These are the exact readout electronics used on the COSI-Balloon payload as described in Beechert et al. [2]. Events are then binned into energy spectra for analysis. To calibrate the detector and measure the spectral resolution prior to any radiation or annealing, we collected photopeak events from gamma-ray sources across a range of energies in the MeV band.
In this work, we focus on spectral measurements of the 0.662 MeV line from Cs-137. Due to non-gaussian and bimodal line profiles resulting from irradiation and temperature cycling the detector, we chose to characterize the detector resolution with an effective FWHM measurement. This FWHM is defined as in terms of the standard deviation () of counts in the Cs-137 peak which traditionally measures the variation of data points about their mean for gaussian and non-gaussian distributions alike. The choice to report resolution results in terms of FWHM rather than is for comparison with spectral resolution FWHM measurements in other works. We recognize that this non-standard FWHM does not permit direct comparison to traditional FWHM measurements of non-gaussian distributions; however, it is well-defined and more easily measured than the traditional FWHM measurements for non-gaussian distributions. This FWHM also has the advantage of converging with the traditional FWHM measurements for gaussian distributions with low background. The robustness of this measurement enables us to consistently track the spectral resolution across the variety of line profiles measured in this work. To determine the FWHM, we constrain peak counts and attempt to remove background counts from each of our distributions by considering energies higher than the energy at which three consecutive bins have counts below 15% of the peak height and lower than 0.670 MeV. We make an exception, fixing a lower bound at 0.475 MeV, for data collected directly after room temperature cycling because the photopeak does not drop off to background levels and instead merges with the Compton backscatter continuum. Each distribution defining the FWHM for hole-collecting strips along with corresponding standard deviation () measurements are shown in Figure 4.
@@ 3.1 Radiation Damage
In order to replicate the proton fluence we might expect over COSI’s mission, we used the proton synchrotron at the James M Slater MD Proton Treatment and Research Center at Loma Linda University Medical Center. We irradiated our detector in two rounds with protons with an average energy of 150 MeV incident on the HV face of the detector. During the first round, the detector received a fluence of p+/cm2. Then, after measuring the spectral resolution of the detector, we delivered another p+/cm2, resulting in a net fluence of p+/cm2. During irradiation, the detector was unbiased and kept under vacuum at operating temperature ( K). Our measurements of the spectral resolution after each round of radiation revealed a significant broadening in photopeaks of hole-collecting strips, characteristic of increased hole trapping due to radiation damage. After completing our post-irradiation measurements, we cycled the detector to room temperature for 336 hours while keeping it under vacuum.
@@ 3.2 Annealing
After each annealing procedure we returned the detector to its operating temperature and collected data from a Cs-137 source, tracking the spectral resolution through the resulting FWHM of the 0.662 MeV line. Radiation damage and annealing at room temperature produced non-gaussian line profiles including low-energy tailing, spectral line shifts, and bimodal distributions. To probe the effectiveness of annealing procedures at different temperatures, we completed anneals at C and C. The two anneals at C were intended to characterize annealing at lower temperatures while, the 5 anneals at C were intended to characterize the progress of the spectral resolution back towards its preradiation value.
@@ 3.3 Numerical Simulations
In parallel to our measurements characterizing the effects of radiation and annealing on the spectral performance, we used numerical simulations to model the photopeak spectrum for a range of trapping products encompassing the anticipated levels of radiation damage. Our numerical charge transport simulations, including the model for charge trapping, are described and tested in Boggs et al. [18].
The measured FWHM of the photopeak has two components: FWHM attributed to the trap density caused by proton damage, and FWHM attributed to a combination of intrinsic trap density, statistical Fano noise, and noise from electronic readout. These components add in quadrature:
For hole-collecting strips, we equated FWHM to the measured preradiation FWHM ( 2.98 keV). We tabulated the FWHM of the Cs-137 photopeak produced by our numerical simulations against their respective trapping products and linearly interpolated between these data points to obtain a relationship between the two parameters. The interpolation and data points are plotted in Figure 1. We used this interpolation to pair the measured FWHM produced at each stage of damage and annealing with the corresponding hole trapping products.
@@ 4 Results
Annealing Temp. | Annealing Time | FWHM | FWHM | FWHM | ||
(∘C) | (Hours) | (keV) | (keV) | (cm) | (keV) | |
Preradiation | – | – | 2.98 | – | 178611Derived in Pike et al. [21] | 4.87 |
Radiation 1 | – | – | 14.4 | 14.1 | 95 | 4.3 |
Radiation 2 | – | – | 38.1 | 38.0 | 36 | 5.0 |
Room Temp | 336 | 115.9 | 115.9 | 10 | 7.1 | |
Anneal 1 | 80 | 24 | 54.5 | 53.9 | 24 | 4.3 |
Anneal 2 | 80 | 48 | 37.2 | 37.1 | 37 | 3.8 |
Anneal 3 | 100 | 24 | 23.3 | 23.1 | 58 | 3.6 |
Anneal 4 | 100 | 48 | 15.7 | 14.4 | 93 | 3.9 |
Anneal 5 | 100 | 96 | 8.41 | 7.86 | 308 | 4.2 |
Anneal 6 | 100 | 192 | 4.39 | 3.22 | 971 | 4.1 |
Anneal 7 | 100 | 192 | 4.08 | 2.79 | 1313 | 4.1 |
@@ 4.1 Preradiation
We measured the FWHM of the Cs-137 photopeak for hole and electron-collecting strips of 2.98 keV and 4.87 keV respectively, prior to any radiation damage or annealing. The broader resolution on the electron-collecting strips is due to HV-coupling and intrinsic electron trapping [22]. The photopeak for hole-collecting strips is pictured in Figure 4.
@@ 4.2 Post Radiation
After the first round of 150 MeV proton irradiation at a fluence of p+/cm2, we measure a peak with a 14.4 keV FWHM for hole-collecting strips – increased nearly by a factor of five from preradiation. After the second round of irradiation, resulting in a net fluence of p+/cm2, our measured FWHM increased to 38.1 keV, over 12 times its preradiation value. In terms of fluence, this reflects an increase of roughly 7.2 keV per p+/cm2 after the first round of radiation and 7.7 keV per p+/cm2 after both rounds of radiation. These measurements demonstrate a roughly linear relationship between detector resolution and net proton fluence. This trend is confirmed in Fourches et al. [13] and Fourches et al. [14] for p-type GeDs, and in Kandel et al. [15] for SPI’s n-type GeDs, all of which underwent neutron irradiation. Likewise, this is seen in Pehl et al. [16] for p-type GeDs which underwent proton irradiation. We observe widened photopeaks and low-energy tailing (Figure 5) characteristic of radiation damage.